| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  214  
 | 
Fri Jul  2 16:14:17 2021 | 
Yuri | Accelerator | Scrapers 4 72Br | Starting 15:15
 
The scrapers are moved very much close to 
each other to eliminate all background particles | 
    | 
| 
  213  
 | 
Fri Jul  2 15:03:37 2021 | 
Pavlos, Alexandre | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000543
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002655
 | 
   | 
| 
  212  
 | 
Fri Jul  2 14:18:04 2021 | 
Pavlos, Alexandre | Runs | Change of the measurment to 72Br | Both resonators will now be retuned to look 
for the isomeric state in 72Br.
 
 | 
   | 
| 
  211  
 | 
Fri Jul  2 13:44:50 2021 | 
Sergey, et al | Accelerator | El cooler | Up to now, attempts to cool the fat line between 
Se and Ar were conducted. We give up it now 
and remove cooling | 
   | 
| 
  210  
 | 
Fri Jul  2 13:37:00 2021 | 
Pavlos, Alexandre | DAQ | Change of the measurement time from 10s to 20s | Last files with the measurement time of 10s 
for the two resonators are:
 
 | 
   | 
| 
  209  
 | 
Fri Jul  2 13:10:00 2021 | 
Yuri | Accelerator | Scrapers | 12:46 moved the TOF detector scraper back 
to -11 mm | 
   | 
| 
  208  
 | 
Fri Jul  2 12:59:21 2021 | 
Pavlos, Alexandre | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000515
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002516
 | 
   | 
| 
  207  
 | 
Fri Jul  2 12:42:54 2021 | 
Gregor | Analysis | Spectra 6:00-12:00 | 0.1s time slices for 245MHz and 410MHz, with 
e143_analyzer.py script
 
 | 
 12x   | 
| 
  206  
 | 
Fri Jul  2 12:18:35 2021 | 
Pavlos, Alexandre | Accelerator | e-cooling | Sergey Litvinov will now try to setup the 
e-cooler to cool the beam in the ESR. 
 
The data taking has been stopped (except | 
   | 
| 
  205  
 | 
Fri Jul  2 11:44:25 2021 | 
Helmut, Yuri | Accelerator | move scrapers to see effect on unknown lines | We close the arc inside and straight section 
inside scrapers to narrow and cut the 70Se 
line.
 | 
   | 
| 
  204  
 | 
Fri Jul  2 10:47:56 2021 | 
Pavlos, Alexandre | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000485
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002370
 | 
   | 
| 
  203  
 | 
Fri Jul  2 08:54:12 2021 | 
Pavlos, Alexandre | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000459
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002242
 | 
   | 
| 
  202  
 | 
Fri Jul  2 07:31:44 2021 | 
The remote Canadian eager beavers (Iris and Guy) | Analysis | 00h to 05h Summed Spectra | Using the scripts I posted previously, I summed 
5.5 hours of data from midnight to 5:30 July 
2nd CEST. 
 | 
       | 
| 
  201  
 | 
Fri Jul  2 06:52:11 2021 | 
The remote Canadian eager beavers (Iris and Guy) | Analysis | Summary of spectra 0:00-6:00 (410 MHz) | Same as before but for 0-6h, 25ms time slices 
starting at 1.04s.
 
 | 
    | 
| 
  200  
 | 
Fri Jul  2 06:51:45 2021 | 
Pavlos, Kanika | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000431
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002106
 | 
   | 
| 
  199  
 | 
Fri Jul  2 06:33:48 2021 | 
The remote Canadian beavers (Iris and Guy) | Analysis | Summary of spectra 4:00-6:00 (410 MHz) | Same as before for 4-5h and 5-6h.
 
No isomer, I think :-( (Sorry, Wolfram!)
 
 | 
    | 
| 
  198  
 | 
Fri Jul  2 05:08:16 2021 | 
Guy Leckenby (Canadian beaver team) | Analysis | e143_analyser.py Summing Script | I've written two scripts that others might 
find useful for the on-line analysis. 
 
  1. The first is a simple bash script that | 
      | 
| 
  197  
 | 
Fri Jul  2 04:58:39 2021 | 
IK, Kanika | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000405
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0001978
 | 
   | 
| 
  196  
 | 
Fri Jul  2 04:27:38 2021 | 
The remote Canadian beavers (Iris and Guy) | Analysis | Summary of spectra 0:00-4:00 (410 MHz) | Same as before for 0-4h on July 2nd.
 
 
> Same as before, for 0-1h on July 2nd
 | 
    | 
| 
  195  
 | 
Fri Jul  2 03:19:49 2021 | 
IK, KK | Runs | drop of intensity | 
 
And the intensity is back again to 5E9 at 
03:20 | 
   |